Consider the following statistics for a small economy in 2008, shown in the table below. What is the employment-to-population ratio?

2008

Working-Age Population

210

Employed

103

Unemployed

23
Answer #1

The employment-to-population ratio is 49.05%

(Employment-to-population ratio = (Employed/Working-Age Population)*100 = (103/210)*100 = 49.05%)